# Serving Large Language Models on Huawei CloudMatrix384

### v2

@@ Abstract
To address these challenges, this paper introduces Huawei CloudMatrix, a next-generation AI datacenter architecture that embodies Huawei’s vision for reshaping the foundation of AI infrastructure. Huawei CloudMatrix384 represents the first production-grade realization of this vision. It integrates 384 Ascend 910C NPUs, 192 Kunpeng CPUs, and other hardware components into a unified supernode, interconnected via an ultra-high-bandwidth, low-latency Unified Bus (UB) network. Unlike conventional hierarchical designs, this architecture enables direct all-to-all communication via UB, allowing compute, memory, and network resources to be dynamically pooled, uniformly accessed, and independently scaled. These architectural features are particularly beneficial for communication-intensive operations such as large-scale MoE expert parallelism and distributed key-value (KV) cache access, making CloudMatrix384 a scalable and high-performance foundation for next-generation LLM serving.
@@
Our extensive evaluation with the DeepSeek-R1 model shows that CloudMatrix-Infer achieves state-of-the-art efficiency without sacrificing accuracy. CloudMatrix-Infer delivers a prefill throughput of 6,688 tokens/s per NPU, and a decode throughput of 1,943 tokens/s per NPU (at ¡50 ms TPOT). These results correspond to compute efficiencies of 4.45 tokens/s/TFLOPS for prefill and 1.29 tokens/s/TFLOPS for decode, both exceeding published results for SGLang on NVIDIA H100 and DeepSeek on NVIDIA H800. CloudMatrix-Infer also effectively manages the throughput-latency trade-off, sustaining a decode throughput of 538 tokens/s per NPU even under the stricter sub-15 ms TPOT constraint. Furthermore, the INT8 quantization on Ascend 910C maintains model accuracy comparable to the official DeepSeek-R1 API across 16 distinct benchmarks.
@@ 1. Introduction
In response to these needs, we present Huawei CloudMatrix, a next-generation AI datacenter architecture built on the principle of fully peer-to-peer high-bandwidth interconnectivity and fine-grained resource disaggregation. We specifically highlight CloudMatrix384, the first production-grade implementation of this innovative architectural concept. CloudMatrix384 is an AI supernode purpose-built for large-scale AI workloads, featuring a fully peer-to-peer interconnected hardware design. It comprises 384 Ascend 910C NPUs and 192 Kunpeng CPUs, interconnected via an ultra-high-bandwidth and low-latency network named unified bus (UB). In particular, this UB network enables direct all-to-all data exchange across all compute and memory components. Unlike conventional hierarchical architectures with uneven intra-node and inter-node interconnect bandwidth, CloudMatrix384 allows the entire supernode to operate as a logically unified, tightly coupled compute entity, embodying the fully peer-to-peer principle that “everything can be pooled, treated equally, and combined freely”. These architectural features are particularly beneficial for communication-intensive operations such as large-scale MoE expert parallelism and distributed KV cache access, making CloudMatrix384 a scalable and high-performance foundation for next-generation LLM serving.
@@ 3.1. Vision for Huawei CloudMatrix
Table 1. Comparison of intra-node and inter-node communication bandwidth and latency in a CloudMatrix384. NPU-related metrics are per-die (each Ascend 910C has two dies). The Ratio column is calculated as Inter-node value / Intra-node value. All reported network bandwidth values denote unidirectional bandwidth.
Unidirectional Bandwidth (GB/s) Latency (µs, 512 Bytes)
Path Operation Inter-node Intra-node Ratio Inter-node Intra-node Ratio
NPU-NPU Read 164 167 0.98 1.9 1.2 1.58
Write 135 137 0.99 2.1 1.3 1.62
NPU-CPU Read 147 151 0.97 1.7 1.0 1.70
Write 107 110 0.97 1.9 1.1 1.73
@@ 3.2. CloudMatrix384 Overview: A Fully Peer-to-Peer Hardware Architecture
CloudMatrix384 is engineered as an AI supernode that integrates 384 Ascend 910C neural-network processing units (NPUs) and 192 Kunpeng central processing units (CPUs), as illustrated in Figure 2. A defining feature of CloudMatrix384 is its peer-to-peer, fully interconnected, ultra-high-bandwidth network that links all NPUs and CPUs via the UB protocol. CloudMatrix384’s UB design is a precursor to the UB-Mesh proposed in (Liao et al., 2025). Each of the 384 NPUs and 192 CPUs connects through UB switches, enabling inter-node communication performance that closely approximates intra-node levels. As shown in Table 1, inter-node bandwidth degradation is under 3%, and inter-node latency increase is less than 1 µs. Given that modern AI workloads are predominantly bandwidth-intensive rather than latency-sensitive, this marginal latency overhead has a negligible impact on the end-to-end performance of AI tasks. Overall, this design allows CloudMatrix384 to function as a tightly-coupled, large-scale logical node with globally addressable compute and memory, facilitating unified resource pooling and efficient workload orchestration.
@@
[…]
@@ 3.3. Hardware Components
3.3.1. Ascend 910C Chip
At the core of CloudMatrix 384 is the HiSilicon Ascend 910C NPU, Huawei’s 2024‐era flagship AI accelerator that succeeds the original Ascend 910B. The 910C is a dual‐die package: two identical compute dies are co-packaged, sharing eight on-package memory stacks and connected by a high-bandwidth cross-die fabric, as shown in Figure 3.
Figure 3. Logical overview of the Huawei Ascend 910C chip, highlighting its dual-die architecture. All reported network bandwidth values denote unidirectional bandwidth.
[…]
3.3.2. Ascend 910C Node
[…]
@@ 3.3.2. Ascend 910C Node
Figure 4. Logical overview of an Ascend 910C node within the CloudMatrix384. All reported network bandwidth values denote unidirectional bandwidth.
@@ 3.3.3. UB Switch System
The CloudMatrix384 supernode spans 16 racks: 12 compute racks, which collectively host the 48 Ascend 910C nodes (384 NPUs in total), and 4 communication racks. These communication racks house the second-tier (L2) UB switches that interconnect all the nodes within the supernode.
Figure 5 illustrates the topology between the on-board first-tier (L1) UB switches (located inside each Ascend 910C node) and the rack-level L2 UB switches. The network is designed to be non-blocking, meaning there is no bandwidth oversubscription at the L2 switching tier. The L2 switches are partitioned into 7 independent sub-planes. Each sub-plane contains 16 L2 UB switch chips, and each L2 switch chip provides $48\times 28\;\mathrm{GB/s}$ ports.
@@ 3.4.1. CANN for Ascend NPUs
Figure 6. The CANN software stack for Huawei Ascend NPUs.
@@ 3.5.2. Architectural Synergy between CloudMatrix384 and DeepSeek Models
[…]
@@
Quantization for Efficiency: INT8 Support. The Ascend 910C’s support for INT8 computation (as described in §3.3.1) presents a valuable opportunity for optimizing the inference performance of DeepSeek models. Quantifying model weights and activations from higher precision formats (like FP16 or BF16) to INT8 can significantly decrease the model’s memory footprint, reduce computational overhead, and lessen memory bandwidth demands during execution. These benefits can translate into improved throughput and reduced latency.
@@ 4.1. Overview: A Peer-to-Peer Serving Architecture with PDC Disaggregation
Each prefill instance is provisioned with 16 Ascend 910C NPUs (32 dies) on CloudMatrix384 and operates with 32-way expert parallelism (EP32). The expert configuration includes 10 experts per rank: one shared expert, eight router experts, and one redundant router expert to support expert parallelism load balancing (EPLB). To further improve efficiency, we employ a hybrid parallelism strategy for MLA computation and apply a microbatch-based pipeline to overlap communication overheads (§4.3).
Each decode instance is allocated a significantly larger NPU group, typically 160 Ascend 910C NPUs (320 dies), to meet the high throughput and low latency demands of autoregressive generation. This setup corresponds to 320-way expert parallelism (EP320) for the MoE layers. Each rank hosts one expert, with the overall configuration consisting of 32 shared experts, 256 distinct router experts, and 32 redundant router experts to facilitate EPLB. To further accelerate decoding, we introduce optimized Ascend-native operators, a pipelined decoding strategy, and multiple-token prediction support, as detailed in §4.2.
@@ 4.2. Tightly-Coupled Decode with Large-scale Expert Parallelism
To address these challenges, we introduce a set of hardware-aware optimization techniques tailored to the CloudMatrix384. First, we present our fused communication operator design that exploits the UB plane for low-latency, high-throughput MoE execution (§4.2.1). Next, we detail our custom MLA implementation for the Ascend 910C (§4.2.2) and describe a microbatch-based decode pipeline that overlaps two execution streams to hide latency (§4.2.3). Finally, we explain how the CloudMatrix-Infer supports multiple-token prediction (MTP), a feature leveraged by DeepSeek-R1 to improve decode throughput (§4.2.4).
@@ 4.2.2. MLA Optimization
Multi-head latent attention (MLA), introduced by DeepSeek, leverages low-rank compression to reduce the spatial footprint of the KV cache and incorporates weight absorption techniques to lower computational costs. While MLA can be deployed on the CloudMatrix384, directly migrating DeepSeek’s operators to Ascend 910C NPUs exposes several performance bottlenecks:
@@
    KV Cache Format Conversion Overhead: To support high-performance matrix computations, the L1 Cache of the Ascend 910C NPU’s AI cube cores (AICs) optimally stores data in an NZ format (a specialized hybrid row-major and column-major layout, resulting in a combined N-shaped and Z-shaped traversal path). However, the KV cache is typically stored in the NPU’s memory using a standard N-Dimensional (ND) format. Consequently, operator internals often need to explicitly convert KV cache data to the NZ format before AICs can perform matrix calculations. This explicit format conversion consumes memory bandwidth and impacts access efficiency, thereby reducing the effective memory bandwidth available for computation.
@@ 4.3. Resource-Efficient Prefill with Hybrid Parallelism and Microbatching
The prefill phase, responsible for processing the input prompt to generate the initial KV cache, significantly impacts time-to-first-token (TTFT) and system throughput. Given its typically compute-intensive nature, achieving high NPU utilization during prefill is paramount. However, this phase often faces challenges such as load imbalances due to heterogeneous input sequence lengths and communication overheads, particularly in complex architectures like MoE models. To address these issues and maximize efficiency on the CloudMatrix384, we propose three key optimizations in CloudMatrix-Infer. First, we introduce a staged hybrid parallelism scheme for MLA computation that overcomes the inherent inefficiencies of conventional data parallelism (§4.3.1). Second, we present a microbatch-based prefill pipeline that exploits the heterogeneous compute and communication units of the Ascend 910C NPU to maximize latency overlap and reduce contention (§4.3.2). Finally, we present the transfer optimizations between prefill and decode phases to minimize the interference to decoding (§4.3.3).
@@ 4.3.2. Microbatch-Based Prefill Pipeline
However, directly porting this strategy to the Ascend 910C NPU on CloudMatrix384 proves inefficient due to architectural mismatches. The pipeline on H800 typically reserves a subset of its streaming multiprocessors (SMs) for communication tasks, enabling concurrency but reducing available compute resources. In contrast, the Ascend 910C offers a heterogeneous compute fabric, which comprises AICs for matrix operations, AIVs for lightweight computation, and SDMA engines for data movement, enabling finer-grained, role-specific task distribution.
@@ 4.5. INT8 Quantization
To achieve high-throughput, low-latency inference for large-scale MoE models such as DeepSeek-V3/R1 on the Ascend 910C platform, we have designed and implemented a training-free, hierarchical INT8 quantization scheme for model weights and activations. This scheme is engineered to maximize computational efficiency and reduce memory footprint while carefully managing potential accuracy degradation. The core components of our approach are detailed below:

### v3

@@ Abstract
To address these challenges, this paper introduces Huawei CloudMatrix, a next-generation AI datacenter architecture that embodies Huawei’s vision for reshaping the foundation of AI infrastructure. Huawei CloudMatrix384 represents the first production-grade realization of this vision. It integrates 384 Ascend 910 NPUs, 192 Kunpeng CPUs, and other hardware components into a unified supernode, interconnected via an ultra-high-bandwidth, low-latency Unified Bus (UB) network. Unlike conventional hierarchical designs, this architecture enables direct all-to-all communication via UB, allowing compute, memory, and network resources to be dynamically pooled, uniformly accessed, and independently scaled. These architectural features are particularly beneficial for communication-intensive operations such as large-scale MoE expert parallelism and distributed key-value (KV) cache access, making CloudMatrix384 a scalable and high-performance foundation for next-generation LLM serving.
@@
Our extensive evaluation with the DeepSeek-R1 model shows that CloudMatrix-Infer achieves state-of-the-art efficiency without sacrificing accuracy. CloudMatrix-Infer delivers a prefill throughput of 6,688 tokens/s per NPU, and a decode throughput of 1,943 tokens/s per NPU (at ¡50 ms TPOT). These results correspond to compute efficiencies of 4.45 tokens/s/TFLOPS for prefill and 1.29 tokens/s/TFLOPS for decode, both exceeding published results for SGLang on NVIDIA H100 and DeepSeek on NVIDIA H800. CloudMatrix-Infer also effectively manages the throughput-latency trade-off, sustaining a decode throughput of 538 tokens/s per NPU even under the stricter sub-15 ms TPOT constraint. Furthermore, the INT8 quantization on Ascend 910 maintains model accuracy comparable to the official DeepSeek-R1 API across 16 distinct benchmarks.
@@ 1. Introduction
In response to these needs, we present Huawei CloudMatrix, a next-generation AI datacenter architecture built on the principle of fully peer-to-peer high-bandwidth interconnectivity and fine-grained resource disaggregation. We specifically highlight CloudMatrix384, the first production-grade implementation of this innovative architectural concept. CloudMatrix384 is an AI supernode purpose-built for large-scale AI workloads, featuring a fully peer-to-peer interconnected hardware design. It comprises 384 Ascend 910 NPUs and 192 Kunpeng CPUs, interconnected via an ultra-high-bandwidth and low-latency network named unified bus (UB). In particular, this UB network enables direct all-to-all data exchange across all compute and memory components. Unlike conventional hierarchical architectures with uneven intra-node and inter-node interconnect bandwidth, CloudMatrix384 allows the entire supernode to operate as a logically unified, tightly coupled compute entity, embodying the fully peer-to-peer principle that “everything can be pooled, treated equally, and combined freely”. These architectural features are particularly beneficial for communication-intensive operations such as large-scale MoE expert parallelism and distributed KV cache access, making CloudMatrix384 a scalable and high-performance foundation for next-generation LLM serving.
@@ 3.2. CloudMatrix384 Overview: A Fully Peer-to-Peer Hardware Architecture
CloudMatrix384 is engineered as an AI supernode that integrates 384 Ascend 910 neural-network processing units (NPUs) and 192 Kunpeng central processing units (CPUs), as illustrated in Figure 2. A defining feature of CloudMatrix384 is its peer-to-peer, fully interconnected, ultra-high-bandwidth network that links all NPUs and CPUs via the UB protocol. CloudMatrix384’s UB design is a precursor to the UB-Mesh proposed in (Liao et al., 2025). Each of the 384 NPUs and 192 CPUs connects through UB switches, enabling inter-node communication performance that closely approximates intra-node levels. The inter-node bandwidth degradation is under 3%, and inter-node latency increase is less than 1 µs. Given that modern AI workloads are predominantly bandwidth-intensive rather than latency-sensitive, this marginal latency overhead has a negligible impact on the end-to-end performance of AI tasks. Overall, this design allows CloudMatrix384 to function as a tightly-coupled, large-scale logical node with globally addressable compute and memory, facilitating unified resource pooling and efficient workload orchestration.
@@
1) UB Plane. The UB plane forms the primary ultra-high-bandwidth scale-up fabric within the supernode. It directly interconnects all 384 NPUs and 192 CPUs in a non-blocking all-to-all topology. UB enables: (1) efficient implementation of fine-grained parallelism strategies such as TP and EP, unconstrained by node boundaries; (2) fast peer-to-peer access to pooled memory (spanning both CPU and NPU memory), which is crucial for efficiently caching model weights and KV caches.
2) RDMA Plane. The RDMA plane enables scale-out communication across CloudMatrix384 supernodes and external RDMA-compatible systems. It currently adopts RDMA over Converged Ethernet (RoCE) to ensure compatibility with standard RDMA stacks. [^1] An alternative design, RDMA over UB, leverages UB’s native support for remote memory access to form a unified UB domain for both intra- and inter-supernode communication. While this approach offers streamlined semantics and avoids protocol translation overhead, the current implementation opts for RoCE to ensure immediate compatibility with existing RDMA libraries and tooling. NPUs are the sole participants in this plane, isolating RDMA traffic from control and storage operations. Key functions include: (1) high-speed transfer of active KV cache data between prefill and decode NPUs during inference; (2) support for distributed training and inference using RDMA-compliant frameworks; (3) low-latency interconnect across supernodes in multi-cluster deployments.
3) VPC Plane. The virtual private cloud (VPC) plane connects the CloudMatrix384 supernode to the broader datacenter network via high-speed NICs (Huawei’s Qingtian card). It operates over standard Ethernet and IP protocols, optionally augmented with UB-over-Ethernet (UBoE). The VPC plane handles: (1) management and control-plane operations such as deployment, monitoring, and scheduling; (2) access to persistent storage, including the object storage service (OBS), the elastic volume service (EVS), and the scalable file system service (SFS); (3) external service communication from CPU-resident workloads, e.g., databases and user interfaces.
@@ 3.3. Hardware Components
3.3.1. Ascend 910 Chip
Figure 3. Logical overview of the Huawei Ascend 910 chip, highlighting its dual-die architecture. All reported network bandwidth values denote unidirectional bandwidth.
At the core of CloudMatrix 384 is the HiSilicon Ascend 910 NPU, Huawei’s 2024‐era flagship AI accelerator. The Ascend 910 is a dual‐die package: two identical compute dies are co-packaged, sharing eight on-package memory stacks and connected by a high-bandwidth cross-die fabric, as shown in Figure 3. Each die contains 24 AI cube (AIC) cores, optimized for matrix and convolution workloads, and 48 AI vector (AIV) cores for element-wise operations. All compute engines support FP16/BF16 and INT8 data types. The 8-bit quantization can be implemented with INT8 precision, enabling computational efficiency comparable to native FP8 hardware without requiring dedicated FP8 support. The two dies communicate over an on-package interconnect. Each Ascend 910 die interfaces with two distinct network planes. 1) UB Plane: The die integrates seven high-speed transceivers to the scale-up UB plane. 2) RDMA Plane: Separately, each die includes a dedicated interface for the scale-out RDMA plane.
3.3.2. Ascend 910 Node
Each compute node in CloudMatrix384 integrates 8 Ascend 910 NPUs, 4 Kunpeng CPUs, and 7 UB switch chips onboard, as illustrated in Figure 4. The 12 processors (8 NPUs and 4 CPUs) connect to these on-board switches via UB links, creating a single-tier UB plane within the node. Each UB switch chip onboard links to the next switching tier in the supernode fabric. Only NPUs participate in the secondary RDMA plane.
@@ 3.3.2. Ascend 910 Node
Figure 4. Logical overview of an Ascend 910 node within the CloudMatrix384. All reported network bandwidth values denote unidirectional bandwidth.
@@ 3.3.3. UB Switch System
The CloudMatrix384 supernode spans 16 racks: 12 compute racks, which collectively host the 48 Ascend 910 nodes (384 NPUs in total), and 4 communication racks. These communication racks house the second-tier (L2) UB switches that interconnect all the nodes within the supernode.
Figure 5 illustrates the topology between the on-board first-tier (L1) UB switches (located inside each Ascend 910 node) and the rack-level L2 UB switches. The network is designed to be non-blocking, meaning there is no bandwidth oversubscription at the L2 switching tier. The L2 switches are partitioned into 7 independent sub-planes. Each sub-plane contains 16 L2 UB switch chips, and each L2 switch chip provides 48 ports.
@@ 3.4.1. CANN for Ascend NPUs
Figure 6. The CANN software stack for Huawei Ascend NPUs.
@@ 3.5.2. Architectural Synergy between CloudMatrix384 and DeepSeek Models
Memory Capacity and Management: Accommodating Large Models and KV Caches. DeepSeek-R1, with parameter counts approaching 671B, requires vast memory resources for both weights and activations, including attention KV caches. CloudMatrix384 provides a huge amount of total NPU-attached memory, enabling distributed storage of model weights through a combination of tensor, pipeline, and expert parallelism. Beyond model weights, LLMs’ attention mechanisms maintain sizable KV caches, especially under long-context or high-batch workloads. CloudMatrix384’s generous memory footprint supports these scenarios, but efficient partitioning and synchronization of KV caches across NPUs remain essential.
@@
Quantization for Efficiency: INT8 Support. The Ascend 910’s support for INT8 computation (as described in §3.3.1) presents a valuable opportunity for optimizing the inference performance of DeepSeek models. Quantifying model weights and activations from higher precision formats (like FP16 or BF16) to INT8 can significantly decrease the model’s memory footprint, reduce computational overhead, and lessen memory bandwidth demands during execution. These benefits can translate into improved throughput and reduced latency.
@@ 4.1. Overview: A Peer-to-Peer Serving Architecture with PDC Disaggregation
Each prefill instance is provisioned with 16 Ascend 910 NPUs (32 dies) on CloudMatrix384 and operates with 32-way expert parallelism (EP32). The expert configuration includes 10 experts per rank: one shared expert, eight router experts, and one redundant router expert to support expert parallelism load balancing (EPLB). To further improve efficiency, we employ a hybrid parallelism strategy for MLA computation and apply a microbatch-based pipeline to overlap communication overheads (§4.3).
Each decode instance is allocated a significantly larger NPU group, typically 160 Ascend 910 NPUs (320 dies), to meet the high throughput and low latency demands of autoregressive generation. This setup corresponds to 320-way expert parallelism (EP320) for the MoE layers. Each rank hosts one expert, with the overall configuration consisting of 32 shared experts, 256 distinct router experts, and 32 redundant router experts to facilitate EPLB. To further accelerate decoding, we introduce optimized Ascend-native operators, a pipelined decoding strategy, and multiple-token prediction support, as detailed in §4.2.
@@ 4.2. Tightly-Coupled Decode with Large-scale Expert Parallelism
To address these challenges, we introduce a set of hardware-aware optimization techniques tailored to the CloudMatrix384. First, we present our fused communication operator design that exploits the UB plane for low-latency, high-throughput MoE execution (§4.2.1). Next, we detail our custom MLA implementation for the Ascend 910 (§4.2.2) and describe a microbatch-based decode pipeline that overlaps two execution streams to hide latency (§4.2.3). Finally, we explain how the CloudMatrix-Infer supports multiple-token prediction (MTP), a feature leveraged by DeepSeek-R1 to improve decode throughput (§4.2.4).
@@ 4.2.2. MLA Optimization
Multi-head latent attention (MLA), introduced by DeepSeek, leverages low-rank compression to reduce the spatial footprint of the KV cache and incorporates weight absorption techniques to lower computational costs. While MLA can be deployed on the CloudMatrix384, directly migrating DeepSeek’s operators to Ascend 910 NPUs exposes several performance bottlenecks:
@@
    KV Cache Format Conversion Overhead: To support high-performance matrix computations, the L1 Cache of the Ascend 910 NPU’s AI cube cores (AICs) optimally stores data in an NZ format (a specialized hybrid row-major and column-major layout, resulting in a combined N-shaped and Z-shaped traversal path). However, the KV cache is typically stored in the NPU’s memory using a standard N-Dimensional (ND) format. Consequently, operator internals often need to explicitly convert KV cache data to the NZ format before AICs can perform matrix calculations. This explicit format conversion consumes memory bandwidth and impacts access efficiency, thereby reducing the effective memory bandwidth available for computation.
@@ 4.3. Resource-Efficient Prefill with Hybrid Parallelism and Microbatching
The prefill phase, responsible for processing the input prompt to generate the initial KV cache, significantly impacts time-to-first-token (TTFT) and system throughput. Given its typically compute-intensive nature, achieving high NPU utilization during prefill is paramount. However, this phase often faces challenges such as load imbalances due to heterogeneous input sequence lengths and communication overheads, particularly in complex architectures like MoE models. To address these issues and maximize efficiency on the CloudMatrix384, we propose three key optimizations in CloudMatrix-Infer. First, we introduce a staged hybrid parallelism scheme for MLA computation that overcomes the inherent inefficiencies of conventional data parallelism (§4.3.1). Second, we present a microbatch-based prefill pipeline that exploits the heterogeneous compute and communication units of the Ascend 910 NPU to maximize latency overlap and reduce contention (§4.3.2). Finally, we present the transfer optimizations between prefill and decode phases to minimize the interference to decoding (§4.3.3).
@@ 4.3.2. Microbatch-Based Prefill Pipeline
However, directly porting this strategy to the Ascend 910 NPU on CloudMatrix384 proves inefficient due to architectural mismatches. The pipeline on H800 typically reserves a subset of its streaming multiprocessors (SMs) for communication tasks, enabling concurrency but reducing available compute resources. In contrast, the Ascend 910 offers a heterogeneous compute fabric, which comprises AICs for matrix operations, AIVs for lightweight computation, and SDMA engines for data movement, enabling finer-grained, role-specific task distribution.
@@ 4.5. INT8 Quantization
To achieve high-throughput, low-latency inference for large-scale MoE models such as DeepSeek-V3/R1 on the Ascend 910 platform, we have designed and implemented a training-free, hierarchical INT8 quantization scheme for model weights and activations. This scheme is engineered to maximize computational efficiency and reduce memory footprint while carefully managing potential accuracy degradation. The core components of our approach are detailed below: